# Stabilizer Ranks, Barnes Wall Lattices and Magic Monotones

### v1

@@ Abstract
In 2024, Kliuchnikov and Schönnenbeck showed a connection between the Barnes Wall lattices, stabilizer states and Clifford operations. In this work, we study their results and relate them to the problem of lower bounding stabilizer ranks. We show the first quantitative lower bound on stabilizer fidelity as a function of stabilizer ranks, which reproduces the linear-by-log lower bound for $\chi_{\delta}({\mathinner{|{H}\rangle}^{\otimes n}})$, i.e., on the approximate stabilizer rank of $\mathinner{|{H}\rangle}^{\otimes n}$. In fact, we show that the lower bound holds even when the fidelity between the approximation and $\mathinner{|{H}\rangle}^{\otimes n}$ is exponentially small, which is currently the best lower bound in this regime.
Next, we define a new magic monotone for pure states, the Barnes Wall norm, and its corresponding approximate variant. We upper bound these monotones by the $CS$-count of state preparation, and also by the stabilizer ranks. In particular, the upper bound given by the $CS$-count is tight, in the sense that we exhibit states that achieve the bound.
Apart from these results, we give a Fidelity Amplification algorithm, which provides a trade-off between approximation error and the stabilizer rank. As a corollary, it gives us a way to compose approximate stabilizer decompositions into approximate decompositions of their tensor products.
Finally, we provide an alternate, elementary proof of the existence and density of product states with maximal stabilizer ranks, which was first proven by Lovitz and Steffan (2022), where they used results from algebraic geometry.
@@ 1 Introduction
Understanding the boundary between classical and quantum computation is a problem of vital importance in the field of quantum information. One way to approach this problem is via the lens of classical simulation of quantum circuits. In this arena, we attempt to quantify the cost of simulating different families of quantum circuits by giving explicit classical algorithms to simulate them and then study how the efficiency of these algorithms scales with the number of input qubits. The first landmark result in this approach was given by Gottesman and Knill [1], which showed that arbitrary quantum circuits using stabilizer states and measurements can be classically simulated efficiently (see also [2]). While these stabilizer circuits are not universal, they can be made to perform universal quantum computation by supplementing them with special non-stabilizer states called magic states.
@@
In [3], Bravyi, Smith and Smolin described a new algorithm to simulate these universal quantum circuits. They decomposed the magic states as linear combinations of a small number of stabilizer states, and then simulated the stabilizer circuits on this combination instead. The minimum number of such stabilizer states needed for a state $\mathinner{|{\psi}\rangle}$ is referred to as the stabilizer rank of $\mathinner{|{\psi}\rangle}$, denoted as $\chi(\mathinner{|{\psi}\rangle})$. For such simulation algorithms, $\chi({\mathinner{|{\psi}\rangle}})$ is effectively the most important metric of efficiency or computational complexity. In particular, if we are able to efficiently find polynomially sized decompositions for magic states like $\mathinner{|{H}\rangle}^{\otimes n}$, then this has various complexity-theoretic implications such as $BQP=BPP$ and even $P=NP$ [4]. Also, instead of finding an exact decomposition in terms of stabilizers, we can also hope to approximate the magic states with linear combination of stabilizer states instead. The corresponding rank in this case is called the $\delta$-approximate stabilizer rank, denoted as $\chi_{\delta}(\mathinner{|{\psi}\rangle})$.
@@
The problem of finding optimal bounds for $\chi(\mathinner{|{\psi}\rangle})$ and $\chi_{\delta}(\mathinner{|{\psi}\rangle})$ for tensor powers of magic states, termed the stabilizer rank problem, has received a lot of attention, with various works chipping away at both the upper and lower bounds, with most results focussing on the ranks of $\mathinner{|{H}\rangle}^{\otimes n}$ or (equivalently) $\mathinner{|{T}\rangle}^{\otimes n}$. Here, $\mathinner{|{H}\rangle}=\cos\frac{\pi}{8}\mathinner{|{0}\rangle}+\sin\frac{\pi}{8}\mathinner{|{1}\rangle}$, and $\mathinner{|{T}\rangle}=\frac{1}{\sqrt{2}}(\mathinner{|{0}\rangle}+e^{i\pi/4}\mathinner{|{1}\rangle})$, and both of these are equally relevant as they help prepare $T$-gates.
@@
For $\chi(\mathinner{|{H}\rangle}^{\otimes n}$, the first lower bound of $\Omega(\sqrt{n})$ was given by Bravyi, Smith and Smolin in [3], with a corresponding upper bound of $O(2^{n/2})$. In a series of works, [3, 5, 6, 4, 7, 8, 9], these lower and upper bounds were improved to $\Omega(\frac{n^{2}}{\mathrm{polylog}(n)})$ and $O(2^{0.4n})$ respectively.
@@
For $\chi_{\delta}(\mathinner{|{H}\rangle}^{\otimes n})$, the best known lower and upper bounds are $\Omega(\frac{n^{2}}{\mathrm{polylog}(n)})$ [7] and $O(2^{0.23n})$[5] for constant $\delta$.
@@
In order to study the stabilizer rank, various other measures of stabilizer complexities have been introduced, including but not limited to: stabilizer fidelity, stabilizer extent and stabilizer alignment (see [5]), with emphasis on relating these to each other and as well as to stabilizer ranks [10, 5, 11]. In particular, we know that tensor products of magic states, such as $\mathinner{|{H}\rangle}^{\otimes n}$, have exponentially large stabilizer extent, and exponentially small stabilizer fidelity.
@@
In a recent work, Kliuchnikov and Schönnenbeck [12] demonstrated that for $n$ qubit Barnes wall lattices, the automorphism group of the lattice, as well as the set of minimum length vectors, are upto phase exactly the Clifford unitaries and stabilizer states respectively. In our work, we study the problem of bounding stabilizer ranks and other notions of stabilizer complexities from the perspective of these recent results.
@@ 1.1 Overview of the Main Results
In Section 3 we prove the first quantitative lower bound on stabilizer fidelity based on the stabilizer rank. Such a bound was shown to exist by Mehraban and Tahmasbi[10], but the result was only existential and did not give a concrete bound. We do this via an intermediate result between the stabilizer rank and stabilizer extent:
@@
Let $\mathinner{|{\phi}\rangle}$ be a pure state such that $\mathinner{|{\phi}\rangle}=\sum_{j=1}^{k}c_{i}\mathinner{|{s_{j}}\rangle}$, where $\mathinner{|{s_{j}}\rangle}$’s are all linearly independent stabilizer states, $c_{j}\in\mathbb{C}$. Then, we have that $\left\|\mathbf{c}\right\|_{1}\leq\sqrt{e}\cdot 2k^{\frac{k+1}{2}}$.
@@
Let $\mathinner{|{\phi}\rangle}$ be a pure state with stabilizer fidelity $F$. For any state $\mathinner{|{\psi}\rangle}$ with $\chi(\psi)=k$, we have $\frac{|\mathinner{\langle{\phi|\psi}\rangle}|}{\sqrt{F}}\leq\sqrt{e}(2k)^{\frac{k+1}{2}}$. In particular, if $\frac{|\mathinner{\langle{\phi|\psi}\rangle}|}{\sqrt{F}}=2^{\Omega(n)}$, then $k=\Omega\left(\frac{n}{\log n}\right)$.
@@
Our lower bounds are the best in this regime, as none of the other results, including the quadratic lower bound in [7], provide any non trivial lower bounds for this.
@@
Applying this result to $\mathinner{|{0}\rangle}^{\otimes n}$ retrieves the best known approximate stabilizer decomposition for $\mathinner{|{H}\rangle}^{\otimes n}$ with rank $O(2^{0.23n})$ first proved in [6].
@@
This result had earlier been shown by Lovitz and Steffan [8], where they had used techniques from algebraic geometry. In our work, we only use the vector space and metric space structure of subspaces spanned by stabilizer states to obtain the result, making these techniques more amenable to being extended to results about approximate stabilizer ranks. Note that any lower bound on the approximate stabilizer ranks of arbitary product states implies the same lower bound (upto log factors) on $\mathinner{|{H}\rangle}^{\otimes n}$, as we can approximate any product state with $O(n\cdot\mathrm{polylog}(n))$ many $T$-gates.
@@ 2 Preliminaries
In this section we will review some background material. We begin with the recalling some basic definitions from the stabilizer rank literature see [5, 8, 3] for more details.
Definition 10 (Also see [5]).
@@
Definition 11 (Also see[13, 5]).
@@
Definition 12 (Also see [5]).
@@
Fact 14.
It was shown in [5] that $F(\mathinner{|{H}\rangle}^{\otimes n})$ = $\cos^{2n}(\frac{\pi}{8})$.
@@ 2.1 Arithmetic for $\mathbb{Z}[i]$
$\mathbb{Q}(i)$ is a field and $\mathbb{Z}[i]$ is a ring which is norm-euclidean. This essentially means that one can define a variation of Euclid’s algorithm for computing the greatest common divisor in this ring. To see this we briefly review some basic properties of $\mathbb{Z}[i]$ (see [14] for more details). The following properties are simply generalizations of the corresponding elementary properties of $\mathbb{Z}$.
@@
Also, since $|\alpha|^{2}\in\mathbb{Z}$ for all $\alpha\in\mathbb{Z}[i]$, it follows directly from the multiplicativity of ||$\cdot$|| ⋅ | that ±$\pm 1,\pm i$ are the only elements with multiplicative inverses. These we refer to as units.
@@
Note that we can only specify the $\gcd$ upto multiplication by units, in our case means that $\gcd$ is unique only upto multiplication with units ±$\pm 1,\pm i$.
@@
An interesting thing that happens in $\mathbb{Z}[i]$ is that elements which are prime in $\mathbb{Z}$ may not be prime in $\mathbb{Z}[i]$. For example, $2=(1+i)(1-i)$.
Fact 18.
@@
Fact 19 (Unique Factorisation).
@@
Fact 20 (Bezout’s lemma).
@@ 2.2 Background on Barnes Wall Lattices
We now review the basic background on Barnes Wall (BW) lattices based on [12, 15], which the reader should consult for more details.
@@
This can be generalised to the notion of a $\mathbb{Z}[i]$-lattice.
@@
The set which we take $\mathbb{Z}$ or $\mathbb{Z}[i]$ linear combinations of are referred to as the generators. Note, that for any lattice $L$, we can create a sublattice $L^{\prime}\subseteq L$ by choosing some elements of $L$ as generators. This sublattice may even lie in a subspace of $V$. We now define the specific $\mathbb{Z}[i]$ lattices we will work with. Define the state |+~$\mathinner{|{\tilde{+}}\rangle}=\frac{1}{1+i}(\mathinner{|{0}\rangle}+\mathinner{|{1}\rangle})=e^{-\frac{i\pi}{4}}\mathinner{|{+}\rangle}$. For each $\mathbf{x}\in\{0,1\}^{n}$, define $\mathinner{|{\tilde{\mathbf{x}}}\rangle}=\bigotimes_{i=1}^{n}\mathinner{|{\tilde{x}_{i}}\rangle}$, where $\mathinner{|{\tilde{x}_{i}}\rangle}=\mathinner{|{0}\rangle}$ if $x_{i}=0$, and $\mathinner{|{\tilde{+}}\rangle}$ otherwise.
Definition 23 (See [12, 15, 16]).
@@
$BW_{n}=\{\sum_{\mathbf{x}\in\{0,1\}}a_{\mathbf{x}}\mathinner{|{\tilde{\mathbf{x}}}\rangle}:~{}a_{\mathbf{x}}\in\mathbb{Z}[i]\}$
Note that we have defined a scaled down version of the lattice as compared to [12], and the results should hold with the corresponding scaling. We did so to ensure that the stabilizer states we work with have norm exactly 1. Furthermore, the usual presentation of the lattice is in terms of generator matrices, but the formulation here suffices for us. Also, [12] works more generally with Barnes wall lattices defined over a larger class of cyclotomic fields, but for our case we restrict to $\mathbb{Q}(i)$.
@@
Also, from the above, since $\{\mathinner{|{\tilde{\mathbf{x}}}\rangle}\}$ for $\mathbf{x}\in\{0,1\}^{n}$ generate $BW_{n}$, we have that $\{(1+i)^{n}\mathinner{|{\tilde{\mathbf{x}}}\rangle}\}$ generate the lattice $BW_{n}^{*}$.
Fact 25.
@@
Fact 27 (Minimal vectors[12]).
For any vector $v\in BW_{n}$, we have $|v|\geq 1$, with $|v|=1$ iff $v$ is a stabilizer state upto phase.
@@
Theorem 28 (Minkowski’s Theorem for lattices [17]).
@@
Here, $\det L$ stands for the volume of fundamental parallelopiped of the lattice, which equals the determinant of the gram matrix $G$ of the generators, which is defined as $G_{ij}=v_{i}^{\dagger}v_{j}$ for the set of generators $\{v_{j}\}$.

### v2

@@ Abstract
In 2024, Kliuchnikov and Schönnenbeck showed a connection between the Barnes Wall lattices, stabilizer states and Clifford operations. In this work, we study their results and relate them to the problem of lower bounding stabilizer ranks. We show the first quantitative lower bound on stabilizer fidelity as a function of stabilizer ranks, which reproduces the linear-by-log lower bound for $\chi_{\delta}({\mathinner{|{H}\rangle}^{\otimes n}})$, i.e., on the approximate stabilizer rank of $\mathinner{|{H}\rangle}^{\otimes n}$. In fact, we show that the lower bound holds even when the fidelity between the approximation and $\mathinner{|{H}\rangle}^{\otimes n}$ is exponentially small, which is currently the best lower bound in this regime. This is shown via an intermediate result between stabilizer fidelity and stabilizer rank , which also gives us an $\Omega(\log n/\log\log n)$ lower bound on the stabilizer rank of pseudorandom quantum states.
Next, we define a new magic monotone for pure states, the Barnes Wall norm, and its corresponding approximate variant. We upper bound these monotones by the $CS$-count of state preparation and also by the stabilizer ranks. In particular, the upper bound given by the $CS$-count is tight, in the sense that we exhibit states that achieve the bound.
Apart from these results, we give a Fidelity Amplification algorithm, which provides a trade-off between approximation error and the stabilizer rank. As a corollary, it gives us a way to compose approximate stabilizer decompositions into approximate decompositions of their tensor products. In particular this recovers the best known approximation for $\mathinner{|{H}\rangle}^{\otimes n}$}. Furthermore we show that this best known approximation is effectively a Barnes Wall lattice approximation. It can also be shown that the Barnes wall norm of this approximation asymptotically matches the upper bound we present.
Finally, we provide an alternative, elementary proof of the existence and density of product states with maximal stabilizer ranks, which was first proven by Lovitz and Steffan (2022), where they used results from algebraic geometry.
@@ 1 Introduction
Understanding the boundary between classical and quantum computation is a problem of vital importance in the field of quantum information. One way to approach this problem is via the lens of classical simulation of quantum circuits. In this arena, we attempt to quantify the cost of simulating different families of quantum circuits by giving explicit classical algorithms to simulate them and then study how the efficiency of these algorithms scales with the number of input qubits. The first landmark result in this approach was given by Gottesman and Knill [gottesman1998heisenberg], which showed that arbitrary quantum circuits using stabilizer states and measurements can be classically simulated efficiently (see also [Aaronson_2004]). While these stabilizer circuits are not universal, they can be made to perform universal quantum computation by supplementing them with special non-stabilizer states called magic states.
@@
In [smith], Bravyi, Smith and Smolin described a new algorithm to simulate these universal quantum circuits. They decomposed the magic states as linear combinations of a small number of stabilizer states, and then simulated the stabilizer circuits on this combination instead. The minimum number of such stabilizer states needed for a given magic state $\mathinner{|{\psi}\rangle}$ is referred to as the stabilizer rank of $\mathinner{|{\psi}\rangle}$, denoted as $\chi(\mathinner{|{\psi}\rangle})$. For such simulation algorithms, $\chi({\mathinner{|{\psi}\rangle}})$ is effectively the most important metric of efficiency or computational complexity. In particular, if we are able to efficiently find polynomially sized decompositions for magic states like $\mathinner{|{H}\rangle}^{\otimes n}$, then this has various complexity-theoretic implications such as $BQP=BPP$ and even $P=NP$ [Peleg2022lowerbounds]. Also, instead of finding an exact decomposition in terms of stabilizers, we can also hope to approximate the magic states with linear combination of stabilizer states instead. The corresponding rank in this case is called the $\delta$-approximate stabilizer rank, denoted as $\chi_{\delta}(\mathinner{|{\psi}\rangle})$.
@@
The problem of finding optimal bounds for $\chi(\mathinner{|{\psi}\rangle})$ and $\chi_{\delta}(\mathinner{|{\psi}\rangle})$ for tensor powers of magic states, termed the stabilizer rank problem, has received a lot of attention, with various works chipping away at both the upper and lower bounds, with most results focusing on the ranks of $\mathinner{|{H}\rangle}^{\otimes n}$ or (equivalently) $\mathinner{|{T}\rangle}^{\otimes n}$. Here, $\mathinner{|{H}\rangle}=\cos\frac{\pi}{8}\mathinner{|{0}\rangle}+\sin\frac{\pi}{8}\mathinner{|{1}\rangle}$, and $\mathinner{|{T}\rangle}=\frac{1}{\sqrt{2}}(\mathinner{|{0}\rangle}+e^{i\pi/4}\mathinner{|{1}\rangle})$.
@@
For $\chi(\mathinner{|{H}\rangle}^{\otimes n})$, the first lower bound of $\Omega(\sqrt{n})$ was given by Bravyi, Smith and Smolin in [smith], with a corresponding upper bound of $O(2^{n/2})$. In a series of works, [smith, Bravyi2019simulationofquantum, Bravyi_2016, Peleg2022lowerbounds, Saeed, Lovitz2022newtechniques, Qassim2021improvedupperbounds], these lower and upper bounds were improved to $\Omega(\frac{n^{2}}{\mathrm{polylog}(n)})$ and $O(2^{0.4n})$ respectively.
@@
For $\chi_{\delta}(\mathinner{|{H}\rangle}^{\otimes n})$, the best known lower and upper bounds are $\Omega(\frac{n^{2}}{\mathrm{polylog}(n)})$ [Saeed] and $O(2^{0.23n})$[Bravyi2019simulationofquantum] for constant $\delta$.
@@
In order to study the stabilizer rank, various other measures of stabilizer complexities have been introduced, including but not limited to: stabilizer fidelity, stabilizer extent and stabilizer alignment (see [Bravyi2019simulationofquantum]), with emphasis on relating these to each other and as well as to stabilizer ranks [mehraban2024improved, Bravyi2019simulationofquantum, Heimendahl2021stabilizerextentis]. In particular, we know that tensor products of magic states, such as $\mathinner{|{H}\rangle}^{\otimes n}$, have exponentially large stabilizer extent, and exponentially small stabilizer fidelity.
@@
In a recent work, Kliuchnikov and Schönnenbeck [kliuchnikov2024stabilizer] demonstrated that for $n$ qubit Barnes Wall lattices, the automorphism group of the lattice, as well as the set of minimum length vectors, are upto phase exactly the Clifford unitaries and stabilizer states respectively. In our work, we study the problem of bounding stabilizer ranks and other notions of stabilizer complexities from the perspective of these recent results.
@@ 1.1 Overview of the Main Results
In Section 3 we prove the first quantitative lower bound on stabilizer fidelity based on the stabilizer rank. Such a bound was shown to exist by Mehraban and Tahmasbi[mehraban2024improved], but the result was only existential and did not give a concrete bound. We do this via an intermediate result between the stabilizer rank and stabilizer extent:
@@
Let $\mathinner{|{\phi}\rangle}$ be a pure state such that $\mathinner{|{\phi}\rangle}=\sum_{j=1}^{k}c_{i}\mathinner{|{s_{j}}\rangle}$, where $\mathinner{|{s_{j}}\rangle}$’s are all linearly independent stabilizer states, $c_{j}\in\mathbb{C}$. Then, we have that $\left\|\mathbf{c}\right\|_{1}\leq\sqrt{e}\cdot(2k)^{\frac{2k+1}{2}}$.
@@
Let $\mathinner{|{\phi}\rangle}$ be a pure state with stabilizer fidelity $F$. For any state $\mathinner{|{\psi}\rangle}$ with $\chi(\psi)=k$, we have $\frac{|\mathinner{\langle{\phi|\psi}\rangle}|}{\sqrt{F}}\leq\sqrt{e}(2k)^{\frac{2k+1}{2}}$. In particular, if $\frac{|\mathinner{\langle{\phi|\psi}\rangle}|}{\sqrt{F}}=2^{\Omega(n)}$, then $k=\Omega\left(\frac{n}{\log n}\right)$.
@@
Our lower bounds are the best in this regime, as none of the other results, including the quadratic lower bound in [Saeed], provide any non trivial lower bounds for this. Moreover, as observed in [mehraban2024improved], such a relation between stabilizer rank and fidelity helps imply a lower bound on stabilizer ranks for pseudorandom states using Theorem 1 from [Grewal2022].
Corollary 4.
Given polynomially many copies of $\mathinner{|{\psi}\rangle}$, we can distinguish, with high probability, the following two cases using a polynomial sized quantum circuit:
  1. 1.
    $\chi(\mathinner{|{\psi}\rangle})=O\left(\frac{\log n}{\log\log n}\right)$
  2. 2.
    $\mathinner{|{\psi}\rangle}$ is a Haar random unit vector.
This improves upon the $O(1)$ lower bound provided in [mehraban2024improved].
@@
    If the measurement does not have uniform probability, then we have the weaker constraint that $\mathcal{N}(\mathinner{|{\phi}\rangle})\geq\mathbb{E}[\mathcal{N}(\mathinner{|{\psi}\rangle})]$, where $\mathinner{|{\psi}\rangle}$ is the random post-measurement state.
@@
Let $\mathinner{|{\phi}\rangle}$ be an $n$-qubit state which can be prepared exactly with $m$ CSCS gates and Clifford operations without any intermediate measurement then:
@@
Applying this result to $\mathinner{|{0}\rangle}^{\otimes n}$ retrieves the best known approximate stabilizer decomposition for $\mathinner{|{H}\rangle}^{\otimes n}$ with rank $O(2^{0.23n})$ first proved in [Bravyi_2016]. In Section 5.1 we further show that the best known approximations for $\mathinner{|{H}\rangle}^{\otimes n}$ are effectively lattice approximations which asymptotically match the bound in Theorem 7.
@@
This result had earlier been shown by Lovitz and Steffan [Lovitz2022newtechniques], where they had used techniques from algebraic geometry. In our work, we only use the vector space and metric space structure of subspaces spanned by stabilizer states to obtain the result, making these techniques more amenable to being extended to results about approximate stabilizer ranks. Note that any lower bound on the approximate stabilizer ranks of arbitary product states implies the same lower bound (upto log factors) on $\mathinner{|{H}\rangle}^{\otimes n}$, as we can approximate any product state with $O(n\cdot\mathrm{polylog}(n))$ many $T$-gates.
@@ 2 Preliminaries
In this section we will review some background material. We begin with the recalling some basic definitions from the stabilizer rank literature see [Bravyi2019simulationofquantum, Lovitz2022newtechniques, smith] for more details.
Definition 11 (Also see [Bravyi2019simulationofquantum]).
@@
Definition 12 (Also see[PhysRevLett.116.250501, Bravyi2019simulationofquantum]).
@@
Definition 13 (Also see [Bravyi2019simulationofquantum]).
@@
Definition 15.
The set of $n$-qubit stabilizer states is $\{C\mathinner{|{0^{n}}\rangle}:\text{ $C$ is an $n$ qubit Clifford Unitary}\}$.
@@
Fact 16.
It was shown in [Bravyi2019simulationofquantum] that $F(\mathinner{|{H}\rangle}^{\otimes n})$ = $\cos^{2n}(\frac{\pi}{8})$.
@@ 2.1 Arithmetic for $\mathbb{Z}[i]$
$\mathbb{Q}(i)$ is a field and $\mathbb{Z}[i]$ is a ring which is norm-euclidean. This essentially means that one can define a variation of Euclid’s algorithm for computing the greatest common divisor in this ring. To see this we briefly review some basic properties of $\mathbb{Z}[i]$ (see [gauss] for more details). The following properties are simply generalizations of the corresponding elementary properties of $\mathbb{Z}$.
@@
Also, since $|\alpha|^{2}\in\mathbb{Z}$ for all $\alpha\in\mathbb{Z}[i]$, it follows directly from the multiplicativity of $|\cdot|$ that $\pm 1,\pm i$ are the only elements with multiplicative inverses. These we refer to as units.
@@
Note that we can only specify the $\gcd$ upto multiplication by units, in our case means that $\gcd$ is unique only upto multiplication with units $\pm 1,\pm i$.
@@
An interesting thing that happens in $\mathbb{Z}[i]$ is that elements which are prime in $\mathbb{Z}$ may not be prime in $\mathbb{Z}[i]$. For example, 2=(1+i)(1i)2=(1+i)(1-i).
Fact 20.
@@
Fact 21 (Unique Factorisation).
@@
Fact 22 (Bezout’s lemma).
@@ 2.2 Background on Barnes Wall Lattices
We now review the basic background on Barnes Wall (BW) lattices based on [kliuchnikov2024stabilizer], which the reader should consult for more details.
@@
A simple example is the lattice $\mathbb{Z}^{n}$ inside $\mathbb{R}^{n}$. This can be generalised to the notion of a $\mathbb{Z}[i]$-lattice.
@@
[…]
Definition 25 (See [kliuchnikov2024stabilizer, Nebe_2002, bwdefn]).
@@
$BW_{n}=\{\sum_{\mathbf{x}\in\{0,1\}^{n}}a_{\mathbf{x}}\mathinner{|{\tilde{\mathbf{x}}}\rangle}:~a_{\mathbf{x}}\in\mathbb{Z}[i]\}$
Note that we have defined a scaled down version of the lattice as compared to [kliuchnikov2024stabilizer], and the results should hold with the corresponding scaling. We did so to ensure that the stabilizer states we work with have norm exactly 1.111In [kliuchnikov2024stabilizer], for $\mathbb{Q}(i)$, the lattice that they work with is created by $\mathbb{Z}[i]$ linear combinations of $(1+i)^{n}\mathinner{|{\tilde{x}}\rangle}$ Furthermore, the usual presentation of the lattice is in terms of generator matrices, but the formulation here suffices for us. Also, [kliuchnikov2024stabilizer] works more generally with Barnes Wall lattices defined over a larger class of cyclotomic fields, but for our case we restrict to $\mathbb{Q}(i)$.
@@
It can easily be seen that the set of elements satisfying the latter condition is indeed closed under addition, however, showing that the set is exactly $BW_{n}$ is a non trivial exercise. Also, from the above, since $\{\mathinner{|{\tilde{\mathbf{x}}}\rangle}\}$ for $\mathbf{x}\in\{0,1\}^{n}$ generate $BW_{n}$, we have that $\{(1+i)^{n}\mathinner{|{\tilde{\mathbf{x}}}\rangle}\}$ generate the lattice $BW_{n}^{*}$.
Fact 27.
@@
Fact 29 (Minimal vectors[kliuchnikov2024stabilizer]).
For any vector $v\in BW_{n}$, we have $\left\|v\right\|\geq 1$, with $\left\|v\right\|=1$ iff $v$ is a stabilizer state upto phase. [^2] In [kliuchnikov2024stabilizer], since they work with a scaled up lattice, the norm of the minimal vectors in that lattice is $2^{n/2}$. Moreover, in [kliuchnikov2024stabilizer], they work with a more general norm of ${\rm{Tr~}}_{K/Q}[\left\|v\right\|_{2}^{2}]$, where $K=\mathbb{Q}(\zeta)\cap\mathbb{R}$, for the case when the lattice is defined over $\mathbb{Z}[\zeta]$. For our case, since we work with $\mathbb{Z}[i]$, we have $K=Q$, and thus the norm reduces to just $\left\|v\right\|^{2}$.
@@
Theorem 30 (Minkowski’s Theorem for lattices [olds2000geometry]).
@@
Here, $\det L$ stands for the volume of fundamental parallelopiped of the lattice, which equals the square root of the determinant of the gram matrix $G$ of the generators, which is defined as $G_{ij}=v_{i}^{\dagger}v_{j}$ for the set of generators $\{v_{j}\}$.